OpenStudy (annalee607):

What are the center and radius of the circle with equation (x+2)^2+(y-9)^2=16

OpenStudy (mathmale):

Compare your "(x+2)^2+(y-9)^2=16" to the standard equation of a circle with center at (h,k) and radius r:\[(x-h)^2+(y-k)^2=r^2\] Can you read off the values of h, k and r through this comparison?

OpenStudy (mathmale):

Compare (y-9) to (y-k). This is the easiest comparison. What is the value of k?

OpenStudy (annalee607):

9

OpenStudy (mathmale):

Yes, and so the center of your circle is (h, 9). What is the x-coordinate of the center? Do another comparison. Be very careful with your signs + and - .

OpenStudy (annalee607):

2

OpenStudy (mathmale):

No. Try again. Compare (x-k) to (x+2). Are the signs the same?

OpenStudy (annalee607):

-2

OpenStudy (mathmale):

Look at the right side of your equation: 16. thios represents r^2. What is r?

OpenStudy (mathmale):

So, the coordinates of the center of your circle are ( ? , ? )

OpenStudy (annalee607):

(-2,9) Radius 4

OpenStudy (mathmale):

Perfect. Review our discussion. Then you'll be done.